OpenStudy (anonymous):

beth is two years older than julio. gerald is twice as old as beth. debra is twice as old as gerald. the sum of their ages is 38. how old is beth?

OpenStudy (anonymous):

j = j b = 2 + j g = 2b d = 2g j + b + g + d = 38 change everything to julio's age j + j + 2 + 2(j + 2) + 2(2(j + 2)) = 38 j + j + 2 + 2j + 4 + 4j + 8 = 38 8j + 14 = 38 8j = 24 j = 3 b = 2 + j beth is 5

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

yup